OpenStudy (anonymous):

How do I symbolize 3*(x+2)^2 + 5 to show 3x^2 + 12x + 27

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (a + b)^2 = a^2 + 2ab + b^2 }\) Hint..

Parth (parthkohli):

Also, remember PEMDAS.

OpenStudy (anonymous):

3(2x^2 + 4x + 4 + 5)...?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3(x + 2)(x + 2) + 5 }\)

OpenStudy (anonymous):

I want to know how to symbolize it correcly.

Parth (parthkohli):

What do you mean by symbolise?

OpenStudy (anonymous):

ParthKohli the 3 in the equation you put only multiplies the (x+2)^2

Parth (parthkohli):

\(\Large \color{MidnightBlue}{ \brack3[(x + 2)(x + 2)] + 3}\)

OpenStudy (anonymous):

I did completing the square but I want to know how I should symbolize the factorization process.

OpenStudy (anonymous):

I will post my thoughts okay then you can tell me if I did it right

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3(x^2 + 4x + 4) + 3 }\)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3x^2 + 12x + 12 + 5 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3x^2 + 12x + 17 }\)

OpenStudy (anonymous):

The question is, complete the square of \[3x ^{2} + 12x + 9\]

Parth (parthkohli):

Okay, that's easy.

Parth (parthkohli):

Go ahead, I'll correct if any mistake

OpenStudy (anonymous):

I know but I want to know how to symbolize it in vertex mode

OpenStudy (anonymous):

My book says do this this 3(x^2 + 4x + 9) ) 3(x^2 + 4x + )+ 9) 3(x^2 + 4x + 4)+9 - 4) 3(x+3)^2 + 5

OpenStudy (anonymous):

My book says to put the 9 outside of a bracket that you make @ParthKohli

OpenStudy (anonymous):

But I think it's better to keep inside so it's mathematically correct

OpenStudy (anonymous):

3x^2 + 12x + 27 /3 /3 /3 3(x^2 + 4x + 9), 3(x^2 + 4x +4 +9 - 4) (x^2 + 4x + 4) = (x+2)^2 3((x + 2)^2 +5)

OpenStudy (anonymous):

Holy crap I figured it out myself!!!!!!!!!!!!!